# Perfect state transfer between real pure states

### v1

@@ Abstract
Pure states correspond to one-dimensional subspaces of $\mathbb{C}^{n}$ represented by unit vectors. In this paper, we develop the theory of perfect state transfer (PST) between real pure states with emphasis on the adjacency and Laplacian matrices as Hamiltonians of a graph representing a quantum spin network. We characterize PST between real pure states based on the spectral information of a graph and prove three fundamental results: (i) every periodic real pure state $\mathbf{x}$ admits perfect state transfer with another real pure state $\mathbf{y}$, (ii) every connected graph admits perfect state transfer between real pure states, and (iii) for any pair of real pure states $\mathbf{x}$ and $\mathbf{y}$ and for any time $\tau$, there exists a real symmetric matrix $M$ such that $\mathbf{x}$ and $\mathbf{y}$ admits perfect state transfer relative to $M$ at time $\tau$. We also determine all real pure states that admit PST in complete graphs, complete bipartite graphs, paths, and cycles. This leads to a complete characterization of pair and plus state transfer in paths and complete bipartite graphs. We give constructions of graphs that admit PST between real pure states. Finally, using results on the spread of graphs, we prove that among all $n$-vertex simple unweighted graphs, the least minimum PST time between real pure states relative to the Laplacian is attained by any join graph, while the it is attained by the join of an empty graph and a complete graph of appropriate sizes relative to the adjacency matrix.
@@ 1 Introduction
where $i^{2}=-1$ and $M$ is a real symmetric matrix called the Hamiltonian associated with $G$. Here, $M$ is indexed by the vertices of $G$ such that $M_{u,v}=0$ if and only if there is no edge between $u$ and $v$. Note that $U(t)$ is a complex, symmetric and unitary matrix for each $t\in\mathbb{R}$. We write $U_{M}(t)$ as $U(t)$ if $M$ is clear from the context. Sometimes, we take $M$ to be $A$ or $L$. But unless otherwise stated, our results apply to any real symmetric matrix $M$ that respects the adjacencies of $G$. We denote the $m\times n$ all-ones matrix and $n\times n$ identity matrix by $J_{m,n}$ and $I_{n}$, respectively. We write these matrices as $J$ and $I$ if the context is clear.
@@
In this paper, we develop the theory of perfect state transfer between real pure states in weighted graphs. In Sections 2 and 4, we extend the concept of eigenvalue supports and strong cospectrality to real pure states. Section 3 deals with periodicity. In particular, we show that periodicity of a real pure state with nonnegative entries is relatively rare whenever $M$ has nonnegative entries, a result that can be viewed as an extension of the relative rarity of vertex periodicity [godsil2010can]. We also provide a simple formula for calculating the minimum period of a periodic real pure state. We devote Section 5 to a characterization of perfect state transfer between real pure states, which extends a characterization of vertex perfect state transfer due to Coutinho [Coutinho2014]. We establish three important facts in this section: (i) every periodic real pure state $\mathbf{x}$ admits perfect state transfer with another real pure state $\mathbf{y}$, (ii) every connected graph admits perfect state transfer between real pure states, and (iii) for any pair of real pure states $\mathbf{x}$ and $\mathbf{y}$ and for any time $\tau$, there exists a real symmetric matrix $M$ such that $\mathbf{x}$ and $\mathbf{y}$ admits perfect state transfer relative to $M$ at time $\tau$. In Sections 6, 7 and 10, we characterize perfect state transfer between real pure states in complete graphs, cycles, paths and complete bipartite graph. As a consequence, we obtain a characterization of pair and plus state transfer in paths and complete bipartite graphs. While only a few cycles and paths admit vertex perfect state transfer, it turns out that there are infinite families of such graphs that admit perfect state transfer between real pure states. Sections 8 and 9 are dedicated to constructions of graphs with PST between real pure states. In Section 11, we utilize results on the spread of graphs to establish that amongst all $n$-vertex unweighted graphs, the minimum PST time between real pure states for the Laplacian case is attained by any join graph, while for the adjacency case, it is attained by the join of an empty graph and a complete graph of appropriate sizes, provided that $n$ is sufficiently large. Finally, in Section 12, we determine closed form expressions for $\frac{d^{k}f}{dt^{k}}\rvert_{\tau}$, where $f(t)=|\mathbf{y}^{\top}U(t)\mathbf{x}|^{2}$ is the fidelity of transfer between real vectors $\mathbf{x}$ and $\mathbf{y}$, and $f(\tau)=1$. We give sharp bounds on $\frac{d^{2}f}{dt^{2}}\rvert_{\tau}$, which describes the sensitivity of the fidelity at time $\tau$ with respect to the readout time.
@@
That is, $\mathbf{y}=\gamma^{2k+1}U(t)\mathbf{x}$. From this, we deduce that for every $\gamma\in\mathbb{C}$, there is at most one real vector in the set $\{\gamma U(t)\mathbf{x}:t>0\}$ distinct from ±$\pm\mathbf{x}$. This demonstrates that the existence perfect state transfer between real pure states is a special occurrence, and therefore warrants an investigation.
@@ 2 Eigenvalue supports
For connected graphs, a vertex state $\mathbf{e}_{u}$ is not an eigenvector for $M$, and hence is not a fixed state.
@@
The covering radius of a set $S\subseteq V(G)$ is the least nonnegative integer $r$ such that each vertex of $G$ is at distance at most $r$ from each vertex in $S$. The covering radius of a vector $\mathbf{x}\in\mathbb{R}^{n}$ is defined to be the covering radius of the set $S=\{u\in V(G):\mathbf{x}^{\top}\mathbf{e}_{u}\neq 0\}$. We state Lemma 4.1 in [godsil2012controllable].
@@ 3 Periodicity
    If $m\geq 3$ and $\mathbf{x}$ is periodic in $G$, then $\rho=\frac{2\pi q}{\lambda_{1}-\lambda_{2}}$, where $q=\operatorname{lcm}(q_{3},\ldots,q_{m})$ and the $p_{j}$s and $q_{j}$s are integers such that $\frac{\lambda_{1}-\lambda_{j}}{\lambda_{1}-\lambda_{2}}=\frac{p_{j}}{q_{j}}$ and $\operatorname{gcd}(p_{j},q_{j})=1$.
@@
By Theorem 8, we may write each $\lambda_{j}=\frac{1}{2}(a+b_{j}\sqrt{\Delta})$, where $\Delta\geq 1$. Thus, each $q_{j}$ in Lemma 10(2) can be written as $q_{j}=\frac{b_{1}-b_{2}}{g_{j}}$, where $g_{j}=\operatorname{gcd}(b_{1}-b_{j},b_{1}-b_{2})$. Therefore, $q=\operatorname{lcm}(q_{3},\ldots,q_{m})=\operatorname{lcm}\left(\frac{b_{1}-b_{2}}{g_{3}},\ldots,\frac{b_{1}-b_{2}}{g_{m}}\right)=\frac{b_{1}-b_{2}}{g}=\frac{\lambda_{1}-\lambda_{2}}{g\sqrt{\Delta}}$. Applying Lemma 10(2) then yields $\rho=\frac{2\pi q}{\lambda_{1}-\lambda_{2}}=\frac{2\pi}{g\sqrt{\Delta}}$. ∎
@@
Let $r$ be the covering radius of $\mathbf{x}$. Since $M$ and $\mathbf{x}$ are entrywise nonnegative, Lemma 5 yields $r+1\leq|\sigma_{\mathbf{x}}(M)|$. Combining this with Lemma 10(1) gives us (1). Now, let $\rho(M)$ denote the spectral radius of $M$. As $|\sigma_{\mathbf{x}}(M)|\geq 3$, $\sigma_{\mathbf{x}}(M)$ is closed under taking algebraic conjugates and $\mathbf{x}$ is periodic, Corollary 9 yields $|\sigma_{\mathbf{x}}(M)|\leq 2\rho(M)+1$. Since $\rho(M)\leq k$, we have $r+1\leq|\sigma_{\mathbf{x}}(M)|\leq 2\rho(M)+1\leq 2k+1,$ which yields the desired result in (2). ∎
@@
Theorem 12 applies to $A$. Further, since we may take $M=kI-L$, Theorem 12 also applies to $L$.
@@
For each $k>0$, there are only finitely many connected graphs $G$ with positive integer weights and maximum degree at most $k$ such that a vector $\mathbf{x}\neq 0$ with nonnegative rational entries is periodic relative to $A$ or $L$.
@@
Assume that $G$ is a connected unweighted graph with maximum degree $k$ that admits PST between vectors $\mathbf{x}$ and $\mathbf{y}$ with nonnegative rational entries. By Lemma 2, $\sigma_{\mathbf{x}}(M)$ is closed under taking algebraic conjugates. Let $r$ be the covering radius of $\mathbf{x}$. Since $\mathbf{x}$ is not fixed, we have $|\sigma_{\mathbf{x}}(M)|\geq 2$, and so applying Theorem 12 to $M\in\{A,kI-L\}$ yields $r\leq 2k$. As $k$ is fixed, there are only finitely many connected unweighted graphs with degree at most $k$ and the covering radius of $\mathbf{x}$ is bounded above by $2k$. This remains true if we assign positive integer weights to $G$. ∎
@@
The argument in the proof of Theorem 14 applies when the Hamiltonian taken is the signless Laplacian matrix. In this case, we obtain the bound $r\leq 4k$ in the above proof in lieu of $r\leq 2k$. Lemma 2 also implies that Theorem 14 applies to entrywise nonnegative vectors $\mathbf{x}\neq{\bf{0}}$ whenever $a\mathbf{x}$ has rational entries for some $a>0$.
@@ 4 Strong cospectrality
The proof of our next result is analogous to that of Theorem 3.1 and Lemma 11.1 in [godsil2017real], and so we omit it here.
@@
    There exists an orthogonal matrix $Q$ that is a polynomial in $M$ and commutes with $M$ such that $Q^{2}=I$ and $Q\mathbf{x}=\mathbf{y}$.
@@
    For some sets $\sigma_{1}$ and $\sigma_{2}$ that partition $\sigma_{\mathbf{x}}(M)$, we have
@@
The proof of the equivalence of (1) and (2) is analogous to that of [godsil2017real, Theorem 11.2]. To prove (1) implies (3), suppose $\mathbf{x}$ and $\mathbf{y}$ are strongly cospectral. Let $\lambda_{k}\in\sigma_{\mathbf{x}}(M)=\sigma_{\mathbf{y}}(M)$. By assumption, $E_{k}\mathbf{x}=\mathbf{u}_{k}$. Thus, $E_{k}\mathbf{x}=E_{k}\mathbf{y}$ if and only if $E_{k}\mathbf{y}=\mathbf{u}_{k}$, while $E_{k}\mathbf{x}=-E_{k}\mathbf{y}$ if and only if $E_{k}\mathbf{y}=-\mathbf{u}_{k}$. Consequently,
@@ 5 Perfect state transfer
      If $\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)=\{\lambda_{1}\}$, then $p_{\ell}$ and $q_{{\ell}}$ are odd for each $\lambda_{\ell}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$.
    2. (b)
      If $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|\geq 2$ with $\lambda_{1},\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, then
@@
    3. (c)
      If $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|\geq 2$ with $\lambda_{1}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$ and $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, then each $q_{j}$ is odd, and $p_{h}$ is even if and only if $\lambda_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$.
@@
To prove (2), suppose PST occurs between $\mathbf{x}$ and $\mathbf{y}$. By Lemma 22(1), $\mathbf{x}$ and $\mathbf{y}$ are strongly cospectral and periodic. By Theorem 8, the latter statement is equivalent to $\frac{\lambda_{1}-\lambda_{j}}{\lambda_{1}-\lambda_{2}}=\frac{p_{j}}{q_{j}}$ for each $j\geq 3$, where $p_{j}$ and $q_{j}$ are integers such that $\operatorname{gcd}(p_{j},q_{j})=1$. In this case, $\tau=\frac{\pi q}{\lambda_{1}-\lambda_{2}}$, where $q=\operatorname{lcm}(q_{3},\ldots,q_{m})$. Now, if $\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)=\{\lambda_{1}\}$, then equation (4) holds if and only if
$\tau\left(\lambda_{1}-\lambda_{\ell}\right)=\pi q\left(\frac{\lambda_{1}-\lambda_{\ell}}{\lambda_{1}-\lambda_{2}}\right)=\frac{\pi qp_{\ell}}{q_{\ell}}\equiv\pi\ \text{(mod $2\pi$)}$
for all $\ell\geq 2$, where $p_{2}=q_{2}=1$. Since $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, we get that $q$ is odd, and so $p_{\ell}$ and $q_{\ell}$ are odd for $\ell\geq 3$. This proves (2a). Now, let $|\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)|\geq 2$ and $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$. Equation (4) holds if and only if
@@
Since $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, we have $\tau\left(\lambda_{1}-\lambda_{2}\right)=\pi q\left(\frac{\lambda_{1}-\lambda_{2}}{\lambda_{1}-\lambda_{2}}\right)=\pi q\equiv 0$ (mod $2\pi$), and so $q$ is even. Now, for equation (5) to hold for each $\lambda_{\ell}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, we need each $\frac{qp_{\ell}}{q_{\ell}}$ to be odd in which case $\nu_{2}(q)=\nu_{2}(q_{\ell})$. Equivalently, the $\nu_{2}(q_{\ell})$’s must all be equal. On the other hand, for (5) to hold for each $\lambda_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, we need each $\frac{qp_{h}}{q_{h}}$ to be even. That is, we must have $\nu_{2}(q)=\nu_{2}(q_{\ell})>\nu_{2}(q_{h})$ for each $\lambda_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$. Thus (2b) holds, and (2c) can be proven similarly. The converse of (2) is straightforward. ∎
@@
In the case that there is PST between $\mathbf{x}$ and $\mathbf{y}$, the minimum PST time is $\frac{\pi}{g\sqrt{\Delta}}$, where $g$ is given in Corollary 11.
@@
If $\mathbf{x}\in\mathbb{R}^{n}\backslash\{{\bf{0}}\}$ is periodic in $G$ at time $\tau$, then $\mathbf{x}$ is involved in perfect state transfer in $G$ at time $\frac{\tau}{2}$.

### v2

@@ Abstract
Pure states correspond to one-dimensional subspaces of $\mathbb{C}^{n}$ represented by unit vectors. In this paper, we develop the theory of perfect state transfer (PST) between real pure states with emphasis on the adjacency and Laplacian matrices as Hamiltonians of a graph representing a quantum spin network. We characterize PST between real pure states based on the spectral information of a graph and prove three fundamental results: (i) every periodic real pure state $\mathbf{x}$ admits perfect state transfer with another real pure state $\mathbf{y}$, (ii) every connected graph admits perfect state transfer between real pure states, and (iii) for any pair of real pure states $\mathbf{x}$ and $\mathbf{y}$ and for any time $\tau$, there exists a real symmetric matrix $M$ such that $\mathbf{x}$ and $\mathbf{y}$ admits perfect state transfer relative to $M$ at time $\tau$. We also determine all real pure states that admit PST in complete graphs, complete bipartite graphs, paths, and cycles. This leads to a complete characterization of pair and plus state transfer in paths and complete bipartite graphs. We give constructions of graphs that admit PST between real pure states. Finally, using results on the spread of graphs, we prove that amongst all $n$-vertex simple unweighted graphs, the least minimum PST time between real pure states relative to the Laplacian is attained by any join graph, while the it is attained by the join of an empty graph and a complete graph of appropriate sizes relative to the adjacency matrix.
Dedicated to Agnes T. Paras on the occassion of her 60th birthday
@@ 1 Introduction
where $i^{2}=-1$ and $M$ is a real symmetric matrix called the Hamiltonian associated with $G$. Here, $M$ is indexed by the vertices of $G$ such that $M_{u,v}=0$ if and only if there is no edge between $u$ and $v$. Note that $U_{M}(t)$ is a complex, symmetric and unitary matrix for each $t\in\mathbb{R}$. We write $U_{M}(t)$ as $U(t)$ if $M$ is clear from the context. Sometimes, we take $M$ to be $A$ or $L$. But unless otherwise stated, our results apply to any real symmetric matrix $M$ that respects the adjacencies of $G$. We denote the $m\times n$ all-ones matrix and $n\times n$ identity matrix by $J_{m,n}$ and $I_{n}$, respectively. We write these matrices as $J$ and $I$ if the context is clear.
@@
In this paper, we develop the theory of perfect state transfer between real pure states in weighted graphs. In Sections 2 and 4, we extend the concept of eigenvalue supports and strong cospectrality to real pure states. Section 3 deals with periodicity. In particular, we show that periodicity of a real pure state with nonnegative entries is relatively rare whenever $M$ has nonnegative rational entries, a result that can be viewed as an extension of the relative rarity of vertex periodicity [godsil2010can]. We also provide a simple formula for calculating the minimum period of a periodic real pure state. We devote Section 5 to a characterization of perfect state transfer between real pure states, which extends a characterization of vertex perfect state transfer due to Coutinho [Coutinho2014]. We establish three important facts in this section: (i) every periodic real pure state $\mathbf{x}$ admits perfect state transfer with another real pure state $\mathbf{y}$, (ii) every connected graph admits perfect state transfer between real pure states, and (iii) for any pair of real pure states $\mathbf{x}$ and $\mathbf{y}$ and for any time $\tau$, there exists a real symmetric matrix $M$ such that $\mathbf{x}$ and $\mathbf{y}$ admit perfect state transfer relative to $M$ at time $\tau$. In Sections 6, 7 and 10, we characterize perfect state transfer between real pure states in complete graphs, cycles, paths and complete bipartite graph. As a consequence, we obtain a characterization of pair and plus state transfer in paths and complete bipartite graphs. While only a few cycles and paths admit vertex perfect state transfer, it turns out that there are infinite families of such graphs that admit perfect state transfer between real pure states. Sections 8 and 9 are dedicated to constructions of graphs with PST between real pure states. In Section 11, we utilize results on the spread of graphs to establish that amongst all $n$-vertex unweighted graphs, the minimum PST time between real pure states for the Laplacian case is attained by any join graph, while for the adjacency case, it is attained by the join of an empty graph and a complete graph of appropriate sizes, provided that $n$ is sufficiently large. Finally, in Section 12, we determine closed form expressions for $\frac{d^{k}f}{dt^{k}}\rvert_{\tau}$, where $f(t)=|\mathbf{y}^{\top}U(t)\mathbf{x}|^{2}$ is the fidelity of transfer between real vectors $\mathbf{x}$ and $\mathbf{y}$, and $f(\tau)=1$. We give sharp bounds on $\frac{d^{2}f}{dt^{2}}\rvert_{\tau}$, which describes the sensitivity of the fidelity at time $\tau$ with respect to the readout time.
@@
That is, $\mathbf{y}=\gamma^{2k+1}U(t)\mathbf{x}$. From this, we deduce that for every $\gamma\in\mathbb{C}$, there is at most one real vector in the set $\{\gamma U(t)\mathbf{x}:t>0\}$ distinct from ±$\pm\mathbf{x}$. This demonstrates that the existence of perfect state transfer between real pure states is a special occurrence, and therefore warrants an investigation.
@@ 2 Eigenvalue supports
If $\mathbf{x}$ is a fixed state, then (4) implies that $U(t)\mathbf{x}=e^{it\lambda}\mathbf{x}$ for any time $t$, where $\lambda$ is an eigenvalue of $M$ with associated eigenvector $\mathbf{x}$. For connected graphs in particular, a vertex state $\mathbf{e}_{u}$ is not an eigenvector for $M$, and hence is not a fixed state.
@@
The covering radius of a set $S\subseteq V(G)$ is the least nonnegative integer $r$ such that each vertex of $G$ is at distance at most $r$ from $S$. The covering radius of a vector $\mathbf{x}\in\mathbb{R}^{n}$ is defined to be the covering radius of the set $S=\{u\in V(G):\mathbf{x}^{\top}\mathbf{e}_{u}\neq 0\}$. We state Lemma 4.1 in [godsil2012controllable].
@@ 3 Periodicity
    If $m\geq 3$ and $\mathbf{x}$ is periodic in $G$, then $\rho=\frac{2\pi q}{\lambda_{1}-\lambda_{2}}$, where $q=\operatorname{lcm}(q_{3},\ldots,q_{m})$ and the $p_{j}$’s and $q_{j}$’s are integers such that $\frac{\lambda_{1}-\lambda_{j}}{\lambda_{1}-\lambda_{2}}=\frac{p_{j}}{q_{j}}$ and $\operatorname{gcd}(p_{j},q_{j})=1$.
@@
By Theorem 8, we may write each $\lambda_{j}=\frac{1}{2}(a+b_{j}\sqrt{\Delta})$, where $\Delta\geq 1$. Thus, each $q_{j}$ in Lemma 10(2) can be written as $q_{j}=\frac{b_{1}-b_{2}}{2g_{j}}$, where $g_{j}=\operatorname{gcd}(b_{1}-b_{j},b_{1}-b_{2})$. Therefore, $q=\operatorname{lcm}(q_{3},\ldots,q_{m})=\operatorname{lcm}\left(\frac{b_{1}-b_{2}}{2g_{3}},\ldots,\frac{b_{1}-b_{2}}{2g_{m}}\right)=\frac{b_{1}-b_{2}}{2g}=\frac{\lambda_{1}-\lambda_{2}}{g\sqrt{\Delta}}$. Applying Lemma 10(2) then yields $\rho=\frac{2\pi q}{\lambda_{1}-\lambda_{2}}=\frac{2\pi}{g\sqrt{\Delta}}$. ∎
@@
Let $r$ be the covering radius of $\mathbf{x}$. Since $M$ and $\mathbf{x}$ are entrywise nonnegative, Lemma 5 yields $r+1\leq|\sigma_{\mathbf{x}}(M)|$. Combining this with Lemma 10(1) gives us (1). Now, let $\rho(M)$ denote the spectral radius of $M$. As $|\sigma_{\mathbf{x}}(M)|\geq 3$, $\sigma_{\mathbf{x}}(M)$ is closed under taking algebraic conjugates and $\mathbf{x}$ is periodic, Corollary 9 yields $|\sigma_{\mathbf{x}}(M)|\leq 2\rho(M)+1$. As $\rho(M)\leq k$, we get $r+1\leq|\sigma_{\mathbf{x}}(M)|\leq 2\rho(M)+1\leq 2k+1,$ which yields the desired result in (2). ∎
@@
Theorem 12 applies to $A$. Since we may take $M=kI-L$, Theorem 12 also applies to $L$.
@@
For each $k>0$, there are only finitely many connected graphs with positive integer weights and maximum degree at most $k$ such that a vector $\mathbf{x}\neq 0$ with nonnegative rational entries is periodic relative to $A$ or $L$.
@@
Assume that $G$ is a connected unweighted graph with maximum degree $k$ and let $\mathbf{x}$ be a vector with nonnegative rational entries that is periodic relative to $M$. By Lemma 2, $\sigma_{\mathbf{x}}(M)$ is closed under taking algebraic conjugates. Let $r$ be the covering radius of $\mathbf{x}$. Since $\mathbf{x}$ is not fixed, we have $|\sigma_{\mathbf{x}}(M)|\geq 2$, and so applying Theorem 12 to $M\in\{A,kI-L\}$ yields $r\leq 2k$. As $k$ is fixed, there are only finitely many connected unweighted graphs with degree at most $k$ and the covering radius of $\mathbf{x}$ is bounded above by $2k$. This remains true if we assign positive integer weights to $G$. ∎
@@
The argument in the proof of Theorem 14 applies when the Hamiltonian taken is the signless Laplacian matrix. In this case, we obtain the bound $r\leq 4k$ in the above proof in lieu of $r\leq 2k$. By Lemma 2, Theorem 14 applies to entrywise nonnegative vectors $\mathbf{x}\neq{\bf{0}}$ whenever $a\mathbf{x}$ has rational entries for some $a>0$.
@@ 4 Strong cospectrality
The proof of our next result is analogous to that of Theorem 3.1 and Lemma 10.1 in [GodsilSmith2024], and so we omit it here.
@@
    There exists an orthogonal matrix $Q$ that is a polynomial in $M$ such that $Q^{2}=I$ and $Q\mathbf{x}=\mathbf{y}$.
@@
    For some nonempty sets $\sigma_{1}$ and $\sigma_{2}$ that partition $\sigma_{\mathbf{x}}(M)$, we have
@@
The proof of the equivalence of (1) and (2) is analogous to that of [GodsilSmith2024, Theorem 10.2]. To prove (1) implies (3), suppose $\mathbf{x}$ and $\mathbf{y}$ are strongly cospectral. Let $\lambda_{k}\in\sigma_{\mathbf{x}}(M)=\sigma_{\mathbf{y}}(M)$. By assumption, $E_{k}\mathbf{x}=\mathbf{u}_{k}$. Thus, $E_{k}\mathbf{x}=E_{k}\mathbf{y}$ if and only if $E_{k}\mathbf{y}=\mathbf{u}_{k}$, while $E_{k}\mathbf{x}=-E_{k}\mathbf{y}$ if and only if $E_{k}\mathbf{y}=-\mathbf{u}_{k}$. Consequently,
@@ 5 Perfect state transfer
      If $\lambda_{1},\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, then
@@
    2. (b)
      If $\lambda_{1}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$ and $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, then each $q_{j}$ is odd, and $p_{h}$ is even if and only if $\lambda_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$.
@@
We now prove (2). Suppose PST occurs between $\mathbf{x}$ and $\mathbf{y}$. By Lemma 22(1), $\mathbf{x}$ and $\mathbf{y}$ are strongly cospectral and periodic. By Theorem 8, the latter statement is equivalent to $\frac{\lambda_{1}-\lambda_{j}}{\lambda_{1}-\lambda_{2}}=\frac{p_{j}}{q_{j}}$ for each $j\geq 3$, where $p_{j}$ and $q_{j}$ are integers such that $\operatorname{gcd}(p_{j},q_{j})=1$. In this case, $\tau=\frac{\pi q}{\lambda_{1}-\lambda_{2}}$, where $q=\operatorname{lcm}(q_{3},\ldots,q_{m})$. To prove (2a), let $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$. Equation (5) holds if and only if
@@
Since $\lambda_{2}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, we have $\tau\left(\lambda_{1}-\lambda_{2}\right)=\pi q\left(\frac{\lambda_{1}-\lambda_{2}}{\lambda_{1}-\lambda_{2}}\right)=\pi q\equiv 0$ (mod $2\pi$), and so $q$ is even. Now, for equation (6) to hold for each $\lambda_{\ell}\in\sigma_{\mathbf{x},\mathbf{y}}^{-}(M)$, we need each $\frac{qp_{\ell}}{q_{\ell}}$ to be odd in which case $\nu_{2}(q)=\nu_{2}(q_{\ell})$. Equivalently, the $\nu_{2}(q_{\ell})$’s must all be equal. On the other hand, for (6) to hold for each $\lambda_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$, we need each $\frac{qp_{h}}{q_{h}}$ to be even. That is, we must have $\nu_{2}(q)=\nu_{2}(q_{\ell})>\nu_{2}(q_{h})$ for each $\lambda_{h}\in\sigma_{\mathbf{x},\mathbf{y}}^{+}(M)$. Thus (2a) holds, and (2b) can be proven similarly. The converse of (2) is straightforward. ∎
@@
In case there is PST between $\mathbf{x}$ and $\mathbf{y}$, the minimum PST time is $\frac{\pi}{g\sqrt{\Delta}}$, where $g$ is given in Corollary 11.
@@
If $\mathbf{x}\in\mathbb{R}^{n}\backslash\{{\bf{0}}\}$ is periodic in $G$ at $\tau$, then $\mathbf{x}$ is involved in perfect state transfer in $G$ at $\frac{\tau}{2}$.